OpenStudy (anonymous):

okay logarithms take two...Express as a single logarithm: loga x - 5loga y

OpenStudy (anonymous):

\[\log_{a}{x} - 5\log_{a}{y} \]

OpenStudy (anonymous):

Property of logarithms: \[a*\log{x} = \log{x^a}\] That means you can switch something between being a power inside and a coefficient outside.

OpenStudy (anonymous):

ok....i remember this...so basically loga x - loga y5..the logarithm basically becomes the exponent right?

OpenStudy (anonymous):

the coefficient become the exponent, but yes. Another property of logarithms: \[\Large \log_a{x} - \log_a{y} = \log_a{\frac{x}{y}}\]

OpenStudy (anonymous):

Basically, if it's subtraction, as long as the base is the same for both logarithms, you can change it into a fraction.

OpenStudy (anonymous):

ok..easy enough..thanks:)